# A Lanczos-Based Algorithmic Approach for Spike Detection in Large Sample Covariance Matrices

### v1

@@ 2.2. Pilot estimation: algorithmic approach for the spiked VASD Stieltjes transform estimation
In our experiments, we typically use the third approach. Theorem 5.5 shows that $\mathrm{O}(\log N)$ Lanczos iterations are sufficient for accurate approximations and give the sufficiency of the first approach, if $C$ is known. As such the maximum number of iteration in Algorithm SR.1 can be set to $\lceil C\log N\rceil$ for $C$ large. In this case, the sequences $\{\widehat{\alpha}_{j}\}$ and $\{\widehat{\beta}_{j}\}$ fluctuate around their deterministic limits with deviations of order $N^{-1/2}$ (see Corollary 5.3), and a natural choice for the convergence tolerances is $\delta_{1}=\delta_{2}=C/\sqrt{N}$. In our numerical examples, setting $q=\lfloor\frac{1}{2}\log N\rfloor$, $\delta=3/\sqrt{N}$, and $C=\lceil\max{6\log N+24,\sqrt{N}}\rceil$ yields reliable results. It is important to note that if a stricter tolerance $\delta$ is initially set, it can be relaxed later without the need to restart the process or incur additional computational costs, as the Cholesky factors have already been computed. We leave an adaptive procedure that incorporates all of these strategies for future work.
@@
1:Input: A lower-bidiagonal semi-infinite matrix (2.3) that is constant after column $n-2$.
@@ 2.3. Averaging procedure and proposed method for spike detection
1:Input: Lower-bidiagonal semi-infinite matrices $\mathcal{L}(j)$, $j=1,2,\ldots,k$, that are constant after column $n(j)-2$, and an integer $q\ll n(j)$.
@@
9:$\widehat{\alpha}\leftarrow\frac{1}{kq}\widehat{\alpha}$$\widehat{\beta}\leftarrow\frac{1}{kq}\widehat{\beta}$
@@
7:Run Algorithm SR.3 on $\mathcal{L}(\mathbf{b}_{j})$ for $j=1,\dots,k$, obtaining ^$\widehat{\mathcal{L}}(\mathbf{b}_{j})$ for $j=1,\dots,k$.
@@
9:     Run Algorithm SR.2 on ^$\widehat{{\mathcal{L}}}(\mathbf{b}_{j})$ to compute2.3 $\widehat{\gamma}_{\pm}$ and ${\widehat{m}}^{j}_{0}(z)$.
@@
With Algorithm P.2 in hand, we are able to develop our final spike detection algorithm. The algorithm works by first running Algorithm P.2 and counting the poles in the computed Stieltjes transforms that lie sufficiently far to the right of the estimate support $[\widehat{\gamma}_{-},\widehat{\gamma}_{+}]$. We refer the reader to Appendix C for an in-depth discussion of two methods (c.f. Algorithms PE.1 and PE.2) to compute these poles.
@@
1:Input: Sample covariance matrix $W$, an integer $k$ and a threshold parameters $C$, $20<\delta<1/$ 2.
@@ 3.1. Orthogonal polynomials, Jacobi matrices and Cholesky factorizations
Assume that $\mu$ is Borel measure on $\mathbb{R}$ with finite mass and compact support. The orthonormal polynomials $(p_{n})_{n\geq 0}$, $p_{n}(\lambda)=p_{n}(\lambda;\mu)$ for $\mu$ are constructed by applying the Gram-Schmidt process to the sequence $(\lambda\mapsto 1,\lambda\mapsto\lambda,\lambda\mapsto\lambda^{2},\ldots)$ with the $L^{2}(\mu)$ inner product [45],
@@
where $\delta_{ij}$ is the Kronecker delta. We also impose that the leading coefficient of $p_{n}(\lambda)$ is positive. If the support of $\mu$ contains at least $N$ points, then $p_{n}(\lambda)$ exists for $10\leq n\leq N-$ 1. The orthonormal polynomials satisfy a symmetric three-term recurrence
@@
One key result is that the Jacobi matrix is connected with the Stieltjes transform of $\mu$ in the sense that $\mathcal{J}(\mu)$ is the unique Jacobi matrix that satisfies [10],
@@
The asymptotic properties of $\mathcal{J}(\mu)$ and its Cholesky factorization $\mathcal{L}(\mu)$ are heavily tied to the regularity conditions outlined in Assumption 1 below, see also [10, 33, 15]. More specifically, if $\mu$ is supported on a single interval with square root behavior at the endpoints, and possibly a few outliers, the corresponding Jacobi operator and its Cholesky factor exhibit simple, exponential asymptotic behavior that can be characterized by the edges of the support.
@@
The result for Jacobi operators in the case $a=-1,b=1$ without discrete contributions was established in [32, 33], and the approach was later extended in [15]. The only piece that requires an extra argument is that [15] required $\{c_{j}\}_{j=1}^{p}$ to be well separated. Yet, the estimates can be seen to hold even if points in this set are allowed to contract at a sufficiently slow exponential rate, see Assumption 1(2). The corresponding argument for the Cholesky factor is provided in [14, Theorem 5.4]. ∎
@@ 3.2. Characterization of Stieltjes transforms using the Cholesky decomposition
Furthermore, $\widehat{m}_{0}(z)$ is the Stieltjes transform of a measure that has its density supported on $[\widehat{\gamma}_{-},\widehat{\gamma}_{+}]$ where
@@ 3.3. Random matrices and fixed point equations for the VASD
where the entries are independent chi random variables with subscripts denoting the degrees of freedom. As $N/M\to c\in(0,1)$, $N\to\infty$, we have, entrywise
@@
We repeat the previous calculation for $\Sigma=\mathrm{diag}(\ell,1,\ldots,1)$ with $\ell>1$. We again apply both stages of Algorithm P.1 explicitly but this time with $\mathbf{b}=\mathbf{e}_{1}$. After applying Algorithm SR.1, it follows that [8],
@@
where the entries are independent chi random variables with subscripts denoting the degrees of freedom. As $N/M\to c\in(0,1)$, $N\to\infty$, we have, entrywise
@@ 4. Large sample covariance matrices and local laws
To simplify our statements, we adopt the concept of stochastic domination [30].
@@ 4.1. The deformed Marchenko-Pastur law and the asymptotics of the outliers
It is well known that [30], in general, the asymptotic density function of the ESD of $\mkern 2.0mu\underline{\mkern-2.0muW\mkern-4.0mu}\mkern 4.0mu_{0}$ follows what is often referred to as the deformed MP law, denoted as $\mu_{\mathrm{dMP}}$, which is best described by its Stieltjes transform. Let $z\in\mathbb{C}_{+}$, the Stieltjes transform $m_{\mathrm{dMP}}(z)$ can be characterized as the unique solution of the equation [30],
@@
We use $\varrho_{\mathrm{dMP}}$ to denote the density associated with $\mu_{\mathrm{dMP}}.$ We work within the framework imposed by the following assumptions, most of which were used in [14].
@@
    On $\Sigma_{0}$ in (1.9): We assume that for some small constant $10<\tau_{1}<$ 1,
@@
Recall that our model of interest (1.1) is formed by adding $r$ spikes to $\Sigma_{0}$, where $r\geq 0$ is some fixed integer. The conditions in Assumption 2(b) rule out the existence of spikes in $\Sigma_{0}$ so that all possible spikes are exclusive to $\Sigma$, and also guarantee that $\varrho_{\mathrm{dMP}}$ displays a regular square-root behavior near the edges $\gamma_{\pm}$. For instance, this condition will be satisfied when the asymptotic spectral distribution of $\Sigma_{0}$ is supported on some interval $[a,b]\subset(0,\infty)$ and its density function is bounded from both above and below; see [30, Example 2.9] or [21, Corollary 3] for more details. Assumption 2(c), for reasons we will now describe, imposes the condition that each $\tilde{\sigma}_{i}$, $1\leq i\leq r$, generate a spike in $W$ that are at an $\mathrm{O}(1)$ distance from $\gamma_{+}$ as $N\to\infty$ [13, 17]. Additionally, Assumption 2(d) imposes a constraint on how closely the outliers of $W$ can approach one another, ensuring that the gaps between them remain at least polynomially small with overwhelming probability.
@@
To be detectable, these spikes in $\Sigma$ should be strong, beyond the so-called BBP transition threshold [4, 42], so that they produce the corresponding outlier eigenvalues in $W$. By Assumption 2, the so-called spiked eigenvalues of $W$ will satisfy
@@ 4.2. Local laws
[…]
@@
We state results relating to the so-called anisotropic local laws [30] for the non-spiked model. Throughout the following, we use the notation $z=\lambda+i\eta$ with $\eta>0$ for the spectral parameter $z$. Fix some small constant $\tau>0$ and consider the set of spectral parameters
@@
In [30], for example, the authors use the alternate definition
@@
for $\delta>0$. From the assumptions made, it follows that $f_{i}(z)$ extends to be uniformly $1/2$-Hölder continuous on $\overline{\mathcal{R}}$. Thus, if the boundary value of $f_{i}(z)$ has a uniform lower bound on $[\gamma_{-},\gamma_{+}]$, by taking $\delta$ sufficiently small, we will have a uniform lower bound on the closure of $\mathcal{R}$. Indeed, Assumption 2(b) implies this lower bound and therefore for a uniform constant $\tau_{2}$ and
@@ 4.3. Asymptotic VESDs (VASDs)
From the concentration of Lipschitz functions on the hypersphere [48, Theorem 5.1.4], there exists an absolute constant $c$ such that if $\mathbf{b}$ is uniformly distributed on $\mathbb{S}^{N-1}$ then
@@ 5. Analysis of the pilot algorithms
Let $\mu$ be a measure supported on a single interval, with a finite number of spikes, and satisfying Assumption 1. Let $\nu$ represent a perturbed (and possibly random) version of $\mu$. The following result from [14] establishes the relation between perturbed and unperturbed Jacobi matrices, as well as their Cholesky factors, asymptotically in terms of the difference of the Stieltjes transform
@@
Let $N$ be a positive integer and suppose $\mu=\mu(N)$ satisfies Assumption 1 for sufficiently large $N$. Suppose further that a measure $\nu=\nu(N)$ is such that $\nu-\sum_{i=1}^{p}w_{j}\delta_{c_{j}},$ has its support inside $\Gamma=\Gamma(\eta)$, where $\Gamma(\eta)$ is the rectangle that is a distance $\eta$ from $[a,b]$ for some $\eta>0$, and assume $\|m(z,\mu-\nu)\|_{L^{\infty}(\Gamma)}\leq E(N,\eta)$. Recall (3.4) and (3.6). If $n\leq C\eta^{-1/2}$, $C>0$ and $\eta=\eta(N)$ is such that $E(N,\eta)\eta^{-1}\to 0$ as $N\to\infty$, then [^6] Note that the factor $\eta^{-1/2}$ in [15, Theorem 2.4] should be $\eta^{-1}$.
@@
It is important to note that $W$ may have eigenvalues exceeding $\gamma_{+}$. Indeed, it is well established in the literature [22] that the largest eigenvalues within the bulk fluctuate around $\gamma_{+}$ on a scale of $N^{-2/3}$ and follow the Tracy-Widom distribution. However, these eigenvalues do not need to be included in $\widecheck{\mu}_{\mathbf{b}}$, as they remain within $\Gamma$ with overwhelming probability. The only spiked eigenvalues accounted for in $\widecheck{\mu}_{\mathbf{b}}$ are the top $r$, which concentrate around the asymptotic spikes $\{\gamma_{j}\}_{j=1}^{k}$. This distinction will play a critical role in Section 6 in identifying the spikes.
@@
and define $m(z)=\mathbf{e}_{1}^{*}(\mathcal{L}\mathcal{L}^{*}-z)^{-1}\mathbf{e}_{1}$ where we emphasize that $\mathcal{L}_{0}$ is constant on the diagonal from the $(n-1)$th entry onwards. Note that
@@
where ˇ$\widecheck{\mathcal{L}}=\mathcal{L}(\widecheck{\mu}_{\mathbf{b}})$. Our argument is structured in two main steps: first, we show that $m(z)$ approximates $\widecheck{m}_{\mathbf{b}}(z)$; then we demonstrate that the estimator $\widehat{m}_{0}(z)$ is relatively close to $m(z)$.
@@
(ˇˇ$\|(\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}-z)^{-1}\|_{2}\leq\frac{1}{\mathrm{Im}\,z}\quad\text{and}\quad\|(\mathcal{L}\mathcal{L}^{*}-z)^{-1}\|_{2}\leq\frac{1}{\mathrm{Im}\,z},$
@@
ˇˇ$\|\mathcal{L}\mathcal{L}^{*}-\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}\|_{2}\leq\|\mathfrak{a}\|_{\infty}+2\|\mathfrak{b}\|_{\infty},$
where $\mathfrak{a}$ and $\mathfrak{b}$ represents the vectors of diagonal and off-diagonal entries of $\mathcal{L}\mathcal{L}^{*}-\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}$, respectively. Using the structure of $\mathcal{L}$ and $\widecheck{\mathcal{L}}$, along with Corollary 5.2, we find that
@@
Using the logarithmic lower bound on $n$, we find ˇˇ$\|\mathcal{L}\mathcal{L}^{*}-\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}\|_{2}=\mathrm{O}_{\prec}(N^{-1/2})$. We conclude that
@@
ˇˇ$\|\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}-\mathcal{L}_{0}\mathcal{L}_{0}^{*}\|_{2}\leq\|\mathfrak{a}_{0}\|_{\infty}+2\|\mathfrak{b}_{0}\|_{\infty},$
where $\mathfrak{a}_{0}$ and $\mathfrak{b}_{0}$ are the vectors of diagonal and off-diagonal entries of ˇˇ$\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}-\mathcal{L}_{0}\mathcal{L}_{0}^{*}$. The vectors $\mathfrak{a}_{0}$ and $\mathfrak{b}_{0}$ can be easily bounded as

### v2

@@ 2.2. Pilot estimation: algorithmic approach for the spiked VASD Stieltjes transform estimation
In our experiments, we typically use the third approach. Theorem 5.5 shows that $\mathrm{O}(\log N)$ Lanczos iterations are sufficient for accurate approximations and give the sufficiency of the first approach, if $C$ is known. As such the maximum number of iteration in Algorithm SR.1 can be set to $\lceil C\log N\rceil$ for $C$ large. In this case, the sequences $\{\widehat{\alpha}_{j}\}$ and $\{\widehat{\beta}_{j}\}$ fluctuate around their deterministic limits with deviations of order $N^{-1/2}$ (see Corollary 5.3), and a natural choice for the convergence tolerances is $\delta_{1}=\delta_{2}=C/\sqrt{N}$. In our numerical examples, setting $q=\lfloor\frac{1}{2}\log N\rfloor$, $\delta=3/\sqrt{N}$, and $C=\lceil\max({6\log N+24,\sqrt{N})}\rceil$ yields reliable results. It is important to note that if a stricter tolerance $\delta$ is initially set, it can be relaxed later without the need to restart the process or incur additional computational costs, as the Cholesky factors have already been computed. We leave an adaptive procedure that incorporates all of these strategies for future work.
@@
1:Input: A lower-bidiagonal semi-infinite matrix (2.3) that is constant after column n2n-2.
@@ 2.3. Averaging procedure and proposed method for spike detection
1:Input: Lower-bidiagonal semi-infinite matrices $\mathcal{L}(j)$, $j=1,2,\ldots,k$, that are constant after column n(j)2n(j)-2, and an integer $q\ll n(j)$.
@@
[…]
@@
7:Run Algorithm SR.3 on $\mathcal{L}(\mathbf{b}_{j})$ for $j=1,\dots,k$, obtaining $\widehat{\mathcal{L}}(\mathbf{b}_{j})$ for $j=1,\dots,k$.
@@
9:  Run Algorithm SR.2 on $\widehat{{\mathcal{L}}}(\mathbf{b}_{j})$ to compute2.3 $\widehat{\gamma}_{\pm}$ and ${\widehat{m}}^{j}_{0}(z)$.
@@
With Algorithm P.2 in hand, we are able to develop our final spike detection algorithm. The algorithm works by first running Algorithm P.2 and counting the poles in the computed Stieltjes transforms that lie sufficiently far to the right of the estimate support $\widehat{\gamma}_{-},\widehat{\gamma}_{+}$]. We refer the reader to Appendix C for an in-depth discussion of two methods (c.f. Algorithms PE.1 and PE.2) to compute these poles.
@@
1:Input: Sample covariance matrix $W$, an integer $k$ and a threshold parameters $C$, $0<\delta<1/2$.
@@ 3.1. Orthogonal polynomials, Jacobi matrices and Cholesky factorizations
Assume that $\mu$ is Borel measure on $\mathbb{R}$ with finite mass and compact support. The orthonormal polynomials $(p_{n})_{n\geq 0}$, $p_{n}(\lambda)=p_{n}(\lambda;\mu)$ for $\mu$ are constructed by applying the Gram-Schmidt process to the sequence $(\lambda\mapsto 1,\lambda\mapsto\lambda,\lambda\mapsto\lambda^{2},\ldots)$ with the $L^{2}(\mu)$ inner product [Szego1939],
@@
where $\delta_{ij}$ is the Kronecker delta. We also impose that the leading coefficient of $p_{n}(\lambda)$ is positive. If the support of $\mu$ contains at least $N$ points, then $p_{n}(\lambda)$ exists for $0\leq n\leq N-1$. The orthonormal polynomials satisfy a symmetric three-term recurrence
@@
One key result is that the Jacobi matrix is connected with the Stieltjes transform of $\mu$ in the sense that $\mathcal{J}(\mu)$ is the unique Jacobi matrix that satisfies [DeiftOrthogonalPolynomials],
@@
The asymptotic properties of $\mathcal{J}(\mu)$ and its Cholesky factorization $\mathcal{L}(\mu)$ are heavily tied to the regularity conditions outlined in Assumption 1 below, see also [DeiftOrthogonalPolynomials, KuijlaarsInterval, Ding2021b]. More specifically, if $\mu$ is supported on a single interval with square root behavior at the endpoints, and possibly a few outliers, the corresponding Jacobi operator and its Cholesky factor exhibit simple, exponential asymptotic behavior that can be characterized by the edges of the support.
@@
The result for Jacobi operators in the case a=1,b=1a=-1,b=1 without discrete contributions was established in [Kuijlaars2003, KuijlaarsInterval], and the approach was later extended in [Ding2021b]. The only piece that requires an extra argument is that [Ding2021b] required $\{c_{j}\}_{j=1}^{p}$ to be well separated. Yet, the estimates can be seen to hold even if points in this set are allowed to contract at a sufficiently slow exponential rate, see Assumption 1(2). The corresponding argument for the Cholesky factor is provided in [Ding2021, Theorem 5.4]. ∎
@@ 3.2. Characterization of Stieltjes transforms using the Cholesky decomposition
Furthermore, $\widehat{m}_{0}(z)$ is the Stieltjes transform of a measure that has its density supported on $\widehat{\gamma}_{-},\widehat{\gamma}_{+}$] where
@@ 3.3. Random matrices and fixed point equations for the VASD
where the entries are independent chi random variables with subscripts denoting the degrees of freedom. As $M\to c\in(0,1), NN\to\infty$, we have, entrywise
@@
We repeat the previous calculation for $\Sigma=\mathrm{diag}(\ell,1,\ldots,1)$ with $\ell>1$. We again apply both stages of Algorithm P.1 explicitly but this time with $\mathbf{b}=\mathbf{e}_{1}$. After applying Algorithm SR.1, it follows that [Bloemendal2013],
@@
where the entries are independent chi random variables with subscripts denoting the degrees of freedom. As $M\to c\in(0,1), NN\to\infty$, we have, entrywise
@@ 4. Large sample covariance matrices and local laws
To simplify our statements, we adopt the concept of stochastic domination [Knowles2017].
@@ 4.1. The deformed Marchenko-Pastur law and the asymptotics of the outliers
It is well known that [Knowles2017], in general, the asymptotic density function of the ESD of $\mkern 2.0mu\underline{\mkern-2.0muW\mkern-4.0mu}\mkern 4.0mu_{0}$ follows what is often referred to as the deformed MP law, denoted as $\mu_{\mathrm{dMP}}$, which is best described by its Stieltjes transform. Let $z\in\mathbb{C}_{+}$, the Stieltjes transform $m_{\mathrm{dMP}}(z)$ can be characterized as the unique solution of the equation [Knowles2017],
@@
We use $\varrho_{\mathrm{dMP}}$ to denote the density associated with $\mu_{\mathrm{dMP}}.$ We work within the framework imposed by the following assumptions, most of which were used in [Ding2021].
@@
    On $\Sigma_{0}$ in (1.9): We assume that for some small constant $0<\tau_{1}<1$,
@@
Recall that our model of interest (1.1) is formed by adding $r$ spikes to $\Sigma_{0}$, where $r\geq 0$ is some fixed integer. The conditions in Assumption 2(b) rule out the existence of spikes in $\Sigma_{0}$ so that all possible spikes are exclusive to $\Sigma$, and also guarantee that $\varrho_{\mathrm{dMP}}$ displays a regular square-root behavior near the edges $\gamma_{\pm}$. For instance, this condition will be satisfied when the asymptotic spectral distribution of $\Sigma_{0}$ is supported on some interval $a,b]\subset(0,\infty)$ and its density function is bounded from both above and below; see [Knowles2017, Example 2.9] or [ElKaroui2007, Corollary 3] for more details. Assumption 2(c), for reasons we will now describe, imposes the condition that each $\tilde{\sigma}_{i}$, $1\leq i\leq r$, generate a spike in $W$ that are at an $\mathrm{O}(1)$ distance from $\gamma_{+}$ as $N\to\infty$ [Ding2021c, Ding2019]. Additionally, Assumption 2(d) imposes a constraint on how closely the outliers of $W$ can approach one another, ensuring that the gaps between them remain at least polynomially small with overwhelming probability.
@@
To be detectable, these spikes in $\Sigma$ should be strong, beyond the so-called BBP transition threshold [Baik2005, perry2018optimality], so that they produce the corresponding outlier eigenvalues in $W$. By Assumption 2, the so-called spiked eigenvalues of $W$ will satisfy
@@ 4.2. Local laws
Let $\widetilde{G}_{1}(z)=(W_{0}-z)^{-1}$, $\widetilde{G}_{2}(z)=(\mkern 2.0mu\underline{\mkern-2.0muW\mkern-4.0mu}\mkern 4.0mu_{0}-z)^{-1}$ and consider
@@
We state results relating to the so-called anisotropic local laws [Knowles2017] for the non-spiked model. Throughout the following, we use the notation $z=\lambda+i\eta$ with $\eta>0$ for the spectral parameter $z$. Fix some small constant $\tau>0$ and consider the set of spectral parameters
@@
In [Knowles2017], for example, the authors use the alternate definition
@@
for $\delta>0$. From the assumptions made, it follows that $f_{i}(z)$ extends to be uniformly $1/2$-Hölder continuous on $\overline{\mathcal{R}}$. Thus, if the boundary value of $f_{i}(z)$ has a uniform lower bound on $\gamma_{-},\gamma_{+}]$, by taking $\delta$ sufficiently small, we will have a uniform lower bound on the closure of $\mathcal{R}$. Indeed, Assumption 2(b) implies this lower bound and therefore for a uniform constant $\tau_{2}$ and
@@ 4.3. Asymptotic VESDs (VASDs)
From the concentration of Lipschitz functions on the hypersphere [Vershynin2018, Theorem 5.1.4], there exists an absolute constant $c$ such that if $\mathbf{b}$ is uniformly distributed on $\mathbb{S}^{N-1}$ then
@@ 5. Analysis of the pilot algorithms
Let $\mu$ be a measure supported on a single interval, with a finite number of spikes, and satisfying Assumption 1. Let $\nu$ represent a perturbed (and possibly random) version of $\mu$. The following result from [Ding2021] establishes the relation between perturbed and unperturbed Jacobi matrices, as well as their Cholesky factors, asymptotically in terms of the difference of the Stieltjes transform
@@
Let $N$ be a positive integer and suppose $\mu=\mu(N)$ satisfies Assumption 1 for sufficiently large $N$. Suppose further that a measure $\nu=\nu(N)$ is such that $\nu-\sum_{i=1}^{p}w_{j}\delta_{c_{j}},$ has its support inside $\Gamma=\Gamma(\eta)$, where $\Gamma(\eta)$ is the rectangle that is a distance $\eta$ from $[a,b]$ for some $\eta>0$, and assume $\|m(z,\mu-\nu)\|_{L^{\infty}(\Gamma)}\leq E(N,\eta)$. Recall (3.4) and (3.6). If $n\leq C\eta^{-1/2}$, $C>0$ and $\eta=\eta(N)$ is such that $E(N,\eta)\eta^{-1}\to 0$ as $N\to\infty$, then [^6] Note that the factor $\eta^{-1/2}$ in [Ding2021b, Theorem 2.4] should be $\eta^{-1}$.
@@
It is important to note that $W$ may have eigenvalues exceeding $\gamma_{+}$. Indeed, it is well established in the literature [Fan2022] that the largest eigenvalues within the bulk fluctuate around $\gamma_{+}$ on a scale of $N^{-2/3}$ and follow the Tracy-Widom distribution. However, these eigenvalues do not need to be included in $\widecheck{\mu}_{\mathbf{b}}$, as they remain within $\Gamma$ with overwhelming probability. The only spiked eigenvalues accounted for in $\widecheck{\mu}_{\mathbf{b}}$ are the top $r$, which concentrate around the asymptotic spikes $\{\gamma_{j}\}_{j=1}^{k}$. This distinction will play a critical role in Section 6 in identifying the spikes.
@@
and define $m(z)=\mathbf{e}_{1}^{*}(\mathcal{L}\mathcal{L}^{*}-z)^{-1}\mathbf{e}_{1}$ where we emphasize that $\mathcal{L}_{0}$ is constant on the diagonal from the (n1)(n-1)th entry onwards. Note that
@@
where $\widecheck{\mathcal{L}}=\mathcal{L}(\widecheck{\mu}_{\mathbf{b}})$. Our argument is structured in two main steps: first, we show that $m(z)$ approximates $\widecheck{m}_{\mathbf{b}}(z)$; then we demonstrate that the estimator $\widehat{m}_{0}(z)$ is relatively close to $m(z)$.
@@
$\|(\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}-z)^{-1}\|_{2}\leq\frac{1}{\mathrm{Im}\,z}\quad\text{and}\quad\|(\mathcal{L}\mathcal{L}^{*}-z)^{-1}\|_{2}\leq\frac{1}{\mathrm{Im}\,z},$
@@
$\|\mathcal{L}\mathcal{L}^{*}-\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}\|_{2}\leq\|\mathfrak{a}\|_{\infty}+2\|\mathfrak{b}\|_{\infty},$
[…]
@@
Using the logarithmic lower bound on $n$, we find $\|\mathcal{L}\mathcal{L}^{*}-\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}\|_{2}=\mathrm{O}_{\prec}(N^{-1/2})$. We conclude that
@@
$\|\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}-\mathcal{L}_{0}\mathcal{L}_{0}^{*}\|_{2}\leq\|\mathfrak{a}_{0}\|_{\infty}+2\|\mathfrak{b}_{0}\|_{\infty},$
where $\mathfrak{a}_{0}$ and $\mathfrak{b}_{0}$ are the vectors of diagonal and off-diagonal entries of $\widecheck{\mathcal{L}}\widecheck{\mathcal{L}}^{*}-\mathcal{L}_{0}\mathcal{L}_{0}^{*}$. The vectors $\mathfrak{a}_{0}$ and $\mathfrak{b}_{0}$ can be easily bounded as